# KKL Observer Synthesis for Nonlinear Systems via Physics-Informed Learning

### v1

@@ 4 Learning Method for KKL Observers
The critical steps of designing KKL observers involve finding the injective map $\mathcal{T}:\mathcal{X}\to\mathcal{Z}$ that satisfies PDE (6), so that the nonlinear system (1) admits a linear up to output injection representation (5), and finding its left-inverse $\mathcal{T}^{*}$, so that a state estimate can be obtained in the original state coordinates. For finding $\mathcal{T}$, one must solve PDE (6), whose explicit solution derived in [4] is given by
$\mathcal{T}(x)=\int_{-\infty}^{0}\exp(A\tau)Bh(\breve{x}(\tau;x))\text{d}\tau$ (10)
@@
However, computing (10) is not practically possible (see [9]) due to the inaccessibility of the backward output map $h(\breve{x}(\tau;x))$ for $\tau<0$ and the infeasibility of computing the integral (10) for every initial point $x\in\mathcal{X}$. Secondly, even if $\mathcal{T}$ is known in some other form [^4] See [8] for some examples. than (10), finding the left-inverse $\mathcal{T}^{*}$ is difficult both analytically and numerically [3]. To avoid these challenges, we propose to instead learn (approximate) the maps $\mathcal{T}$ and $\mathcal{T}^{*}$ using neural networks.
4.1 Problem Definition
Let $\hat{\mathcal{T}}_{\theta}$ and $\hat{\mathcal{T}}^{*}_{\eta}$ be parametrized neural networks that approximate $\mathcal{T}$ and $\mathcal{T}^{*}$, respectively. Here, $\theta\in\mathbb{R}^{n_{\theta}}$ and $\eta\in\mathbb{R}^{n_{\eta}}$ are vectors containing all the weights and biases of the neural networks. Then, the optimal values of $\theta$ and $\eta$ are defined as follows:
[…]
where $x(\tau;\chi)\in\mathcal{X}$ denotes the state at time $\tau\in[0,\infty)$ initialized at $x(0)=\chi\in\mathcal{X}$, and
$\mathcal{P}_{\theta}(x)\coloneq\frac{\partial\hat{\mathcal{T}}_{\theta}}{\partial x}(x)f(x)-A\hat{\mathcal{T}}_{\theta}(x)-Bh(x)$ (12)
is the residual of PDE (6) at $x\in\mathcal{X}$.
4.2 Data Generation
The optimization problems in (11) are mathematically well-defined and can be numerically solved up to local minima by sampling solution trajectories of (1) and (5). Since the matrices $A$ and $B$ are fixed and the function $f$ is known, it is possible to generate trajectories of $x$ and $z$ numerically. The only problem comes from the fact that, given $x_{0}\in\mathcal{X}$, we do not know the initial condition of $z$ because $z_{0}={\mathcal{T}}(x_{0})$. For an arbitrary $z_{0}$, note that the solution of (5) satisfies
[…]
Since $A$ is Hurwitz, $\|\exp(At)z_{0}\|\to 0$ exponentially fast, and the effect of the initial condition $z_{0}$ vanishes from $z(t;z_{0})$ over time $t$. Consequently, for any $\varepsilon>0$ and $t>t_{*}(\varepsilon,z_{0})$ with
$t_{*}(\varepsilon,z_{0})=\frac{1}{\lambda_{\min}(A)}\ln\left(\frac{\varepsilon}{\text{cond}(V)\|z_{0}\|}\right)$
we have
$\left\|\exp(At)z_{0}\right\|$ $\leq\|\exp(At)\|\|z_{0}\|$
$\leq\text{cond}(V)e^{\lambda_{\min}(A)t}\|z_{0}\|$
$\leq\varepsilon$
with $\varepsilon>0$ the upper bound of the impact of the initial condition $z_{0}$ on the norm $\|z(t;z_{0})\|$, $\lambda_{\min}(A)$ the real part of the eigenvalue of $A$ with smallest magnitude, and $V$ obtained from the eigendecomposition of $A=V\Lambda V^{-1}$, which is assumed to be diagonalizable. Therefore,
[…]
@@ 4.2 Data Generation
where $t_{*}\coloneqq t_{*}(\varepsilon,z_{0})$. As time progresses, the trajectory $z(t;z_{0}^{i})$ becomes almost independent of the choice of the initial condition $z_{0}$. This observation leads to the following data generation procedure.
@@
Initial condition sampling: We sample $p\in\mathbb{Z}_{>0}$ initial points $\left\{(x_{0}^{i},z_{0}^{i})\right\}_{i\in[p]}$ uniformly in $\mathcal{X}\times\mathcal{Z}$ using Latin hypercube, orthogonal, or random sampling methods.
@@
Truncation: For all $i\in[p]$, the values of $z$-trajectory for $t_{k}\geq t_{k_{*}}$ are kept and others are discarded, where
@@
The truncation leads to the following datasets:
$X$ =$=\{x(t_{1};x_{0}^{i}),\dots,x(t_{\tau};x_{0}^{i})\}_{i\in[p]}$ (13a)
$Z$ =$=\{z(t_{k_{*}};z_{0}^{i}),\dots,z(t_{\tau};z_{0}^{i})\}_{i\in[p]}.$ (13b)
Note that, for all $i\in[p]$ and $t_{k}\geq t_{k_{*}}$, we have $z(t_{k},z_{0}^{i})\approx\mathcal{T}(x(t_{k},x_{0}^{i}))$.
Figure 1: Overview of our proposed algorithm.
@@ 4.3 Empirical Risk Minimization
The optimization problems in (11) minimize the loss functions that account for the deviation of the neural networks $\hat{\mathcal{T}}_{\theta}$ and $\hat{\mathcal{T}}_{\eta}^{*}$ with respect to the data trajectories. To this end, we can exploit the sampled data
$x^{i}(t_{k})\coloneqq x(t_{k},x_{0}^{i}),\quad z^{i}(t_{k})\coloneqq z(t_{k},z_{0}^{i})$
for learning $\hat{\mathcal{T}}_{\theta}$ and $\hat{\mathcal{T}}_{\eta}^{*}$. We first learn $\hat{\mathcal{T}}_{\theta}$ and then use it to learn $\hat{\mathcal{T}}_{\eta}^{*}$ by approximating the learning problems (11c) and (11d), respectively.
4.3.1 Learning transformation map $\hat{\mathcal{T}}_{\theta}$
The empirical risk associated with (11c) is given by
$\mathcal{L}_{\theta}(X,Z)\coloneqq\frac{1}{p\tau}\sum_{i=1}^{p}\bigg{[}\frac{\tau}{\tau-k_{z}+1}\sum_{k=k_{z}}^{\tau}\|z^{i}(t_{k})-\hat{\mathcal{T}}_{\theta}(x^{i}(t_{k}))\|^{2}\\ +\upsilon\sum_{k=1}^{\tau}\|\mathcal{P}_{\theta}(x^{i}(t_{k}))\|^{2}\bigg{]}$ (14)
where $\upsilon>0$ is a hyperparameter, $X$ and $Z$ are datasets given in (13), and $\mathcal{P}_{\theta}$ is the PDE residual defined in (12). The estimate of $\theta^{\star}$ is given by minimizing the empirical risk $\mathcal{L}_{\theta}$ in (14), i.e.,
$\hat{\theta}^{\star}\in\operatorname*{arg\,min}_{\theta}\mathcal{L}_{\theta}(X,Z).$ (15)
4.3.2 Learning inverse map $\hat{\mathcal{T}}_{\eta}^{*}$
The empirical risk associated with (11d) is given by
$\mathcal{L}_{\eta}(X)\coloneqq\frac{1}{p\tau}\sum_{i=1}^{p}\sum_{k=1}^{\tau}\|x^{i}(t_{k})-\hat{\mathcal{T}}_{\eta}^{*}(\hat{\mathcal{T}}_{\hat{\theta}^{\star}}(x^{i}(t_{k})))\|^{2}$ (16)
where the input to the neural network $\hat{\mathcal{T}}_{\eta}^{*}$ is generated from the neural network $\hat{\mathcal{T}}_{\hat{\theta}^{\star}}$ already learned in (15). The estimate of $\eta^{\star}$ is given by minimizing the empirical risk $\mathcal{L}_{\eta}$ in (16), i.e.,
$\hat{\eta}^{\star}\in\operatorname*{arg\,min}_{\eta}\mathcal{L}_{\eta}(X).$ (17)
Remark 8.
Without the explicit supervision to connect the system’s state space $\mathcal{X}$ to the observer’s state space $\mathcal{Z}$ through PDE residual (12) in empirical risk (14), the neural network $\hat{\mathcal{T}}_{\eta}^{*}$ would only minimize the reconstruction loss $\|x-\hat{\mathcal{T}}_{\eta}^{*}(z)\|$ using the limited number of training points. This makes the model overfit on the training data and hinders the generalization to the unseen data, as is observed in [13]. Moreover, in the extreme case, without adding the PDE, if the neural network $\hat{\mathcal{T}}_{\eta}^{*}$ is rich enough, one could essentially recover the $x$ observation from arbitrary noise $z$. Thus, the learned $\hat{\mathcal{T}}_{\eta}^{*}$ may not necessarily map $\mathcal{T}_{\theta}(x^{\prime})$ close to the point $x^{\prime}$, where $x^{\prime}$ pertains to the test data. $\diamond$
Remark 9.
In our previous work [41], we proposed to simultaneously learn both neural networks $\hat{\mathcal{T}}_{\theta}$ and $\hat{\mathcal{T}}_{\eta}^{*}$. However, it may result in suboptimal parameters $\eta$ because in that case, the empirical risk $\mathcal{L}_{\eta}$ would also depend on $\theta$, which is being simultaneously updated during learning. Therefore, as also reported in [45], the method presented in [41] results in larger approximation (as well as estimation) errors. In the current paper, we exploit the strict hierarchy in the learning problems (11c) and (11d). That is, we observe that learning $\hat{\mathcal{T}}_{\theta}$ is completely decoupled from $\hat{\mathcal{T}}_{\eta}^{*}$. Thus, we propose to first learn $\hat{\mathcal{T}}_{\theta}$ and then use the learned $\hat{\mathcal{T}}_{\hat{\theta}^{\star}}$ to generate $z$-data by passing every $x$-data point through the learned neural network. Once the $\hat{\mathcal{T}}_{\hat{\theta}^{\star}}$ is fixed, we can optimally learn the neural network $\hat{\mathcal{T}}_{\eta}^{*}$ pertaining to the inverse map $\mathcal{T}^{*}$. $\diamond$
4.4 Learning Algorithm
@@
Fig. 1 illustrates the workflow of our proposed learning algorithm. Our procedure deconstructs the autoencoder framework used in previous works such as [44, 41] into learning the transformations $\hat{\mathcal{T}}_{\theta}$ and $\hat{\mathcal{T}}_{\eta}^{*}$ disjointly. This deconstruction serves to address two fundamental issues with the previous design choices:
@@
    The joint encoder-decoder structure is inherently susceptible to conflicting gradients between the different components of the objective, which might result in the optimization getting stuck in bad local minima. This would not only degrade the overall accuracy of the learned solutions but also create the risk of failing to enforce the latent output of the autoencoder to reside within $\mathcal{Z}$, leading to a discrepancy between the input distributions of $\hat{\mathcal{T}}_{\eta}^{*}$ during online operation.

### v2

@@ 5 Learning Method for KKL Observers
Designing KKL observers involves (a) finding the injective map $\mathcal{T}:\mathcal{X}\to\mathcal{Z}$ that satisfies PDE (6), so that the nonlinear system (1) admits a linear up to output injection representation (5), and (b) finding its left-inverse $\mathcal{T}^{*}$, so that a state estimate can be obtained in the original state coordinates. For finding $\mathcal{T}$, one must solve PDE (6), whose explicit solution derived in [4] is given by
$\mathcal{T}(x)=\int_{-\infty}^{0}\exp(A\tau)Bh(\breve{x}(\tau;x))\mathrm{d}\tau$ (10)
@@
However, computing (10) is not practically possible [11] due to the inaccessibility of the backward output map $h(\breve{x}(\tau;x))$ for $\tau<0$ and the infeasibility of computing the integral (10) for every initial point $x\in\mathcal{X}$. Note that, even if $\mathcal{T}$ is known in some other form [^4] See [10] for some examples. than (10), finding the left-inverse $\mathcal{T}^{*}$ is difficult both analytically and numerically [3]. To overcome these challenges, we propose learning the maps $\mathcal{T}$ and $\mathcal{T}^{*}$.
5.1 Learning Problem
We introduce two expectation operators corresponding to the underlying data distributions. Let $\mathcal{D}_{\mathcal{X}}$ be the true data distribution over the compact state space $\mathcal{X}$. The first expectation operator is formally defined with respect to this measure as
$\mathbb{E}_{x\sim\mathcal{D}_{\mathcal{X}}}[\cdot]=\int_{\mathcal{X}}(\cdot)\ \mathrm{d}\mathcal{D}_{\mathcal{X}}(x).$
The injective map $\mathcal{T}:\mathcal{X}\to\mathcal{Z}$ induces a pushforward measure $\mu_{\mathcal{T}}$ on the observer space $\mathcal{Z}$, defined by $\mu_{\mathcal{T}}(Z)=\mathcal{D}_{\mathcal{X}}(\mathcal{T}^{-1}(Z))$ for any measurable set $Z\subset\mathcal{Z}$. The second expectation operator is defined with respect to this measure as
$\mathbb{E}_{z\sim\mu_{\mathcal{T}}}[\cdot]=\int_{\mathcal{Z}}(\cdot)\ \mathrm{d}\mu_{\mathcal{T}}(z).$
Now, we define the risk and error functions associated with the learning of a KKL observer.
Our goal is to find approximations $\hat{\mathcal{T}}_{\theta}\in\mathcal{H}_{\theta}$ and $\hat{\mathcal{T}}_{\eta}^{*}\in\mathcal{H}_{\eta}$ for $\mathcal{T}$ and $\mathcal{T}^{*}$, where $\mathcal{H}_{\theta}$ and $\mathcal{H}_{\eta}$ are the hypothesis classes (e.g., neural network architectures) for the forward and inverse maps. The learned forward map $\hat{\mathcal{T}}_{\theta}$ must satisfy two conditions: (i) it must map $x$ to the correct $z=\mathcal{T}(x)$, and (ii) it must satisfy the governing PDE (6). The forward map error is given by
$R_{\mathcal{T}}(\theta)=\mathbb{E}_{x\sim\mathcal{D}_{\mathcal{X}}}\big[\|\mathcal{T}(x)-\hat{\mathcal{T}}_{\theta}(x)\|^{2}\big]$ (11)
and the PDE error by $R_{\mathrm{PDE}}(\theta)=\mathbb{E}_{x\sim\mathcal{D}_{\mathcal{X}}}\big[\|\mathcal{P}_{\theta}(x)\|^{2}\big]$ where $\mathcal{P}_{\theta}(x)$ is the PDE residual
$\mathcal{P}_{\theta}(x)\coloneq\frac{\partial\hat{\mathcal{T}}_{\theta}}{\partial x}(x)f(x)-A\hat{\mathcal{T}}_{\theta}(x)-Bh(x).$ (12)
We combine these into the physics-informed risk
$R_{1}(\theta)=R_{\mathcal{T}}(\theta)+\nu R_{\mathrm{PDE}}(\theta)$ (13)
where $\nu>0$ is a hyperparameter.
The learned inverse map $\hat{\mathcal{T}}_{\eta}^{*}$ must approximate the true inverse $\mathcal{T}^{*}$ by minimizing the inverse map error
$R_{\mathcal{T}^{*}}(\eta)=\mathbb{E}_{z\sim\mu_{\mathcal{T}}}\big[\|\mathcal{T}^{*}(z)-\hat{\mathcal{T}}_{\eta}^{*}(z)\|^{2}\big]$ (14)
where $\mu_{\mathcal{T}}$ is the pushforward measure on the observer space $\mathcal{Z}$. Moreover, to learn the KKL observer, $\hat{\mathcal{T}}_{\eta}^{*}$ must also reconstruct the state $x$ from the output of the learned forward map $\hat{\mathcal{T}}_{\theta}(x)$. That is, given a fixed, learned map $\hat{\mathcal{T}}_{\theta}$, the learned inverse $\hat{\mathcal{T}}_{\eta}^{*}$ must minimize the reconstruction risk
$R_{2}(\eta\mid\theta)=\mathbb{E}_{x\sim\mathcal{D}_{\mathcal{X}}}\big[\|x-\hat{\mathcal{T}}_{\eta}^{*}(\hat{\mathcal{T}}_{\theta}(x))\|^{2}\big].$ (15)
Our goal is to find parameters $\theta$ and $\eta$ that sequentially minimize $R_{1}(\theta)$ and $R_{2}(\eta\mid\theta)$, respectively. Since the true data distribution $\mathcal{D}_{\mathcal{X}}$ and the true maps $\mathcal{T}$ and $\mathcal{T}^{*}$ are unknown, we minimize empirical approximations of these risks using synthetically generated data.
5.2 Data Generation Procedure
Before defining empirical approximations of these risks, we describe a heuristic for generating synthetic data. Since the observer matrices $A$ and $B$ are assumed to be fixed and the system functions $f$ and $h$ to be known, it is possible to generate trajectories of $x$ and $z$ numerically. However, given $x_{0}\in\mathcal{X}$, one cannot guess the initial condition $z(0)={\mathcal{T}}(x_{0})$ because $\mathcal{T}$ is unknown. Following [44], one could mitigate this because, for an arbitrary $z(0)=z_{0}$, the solution of (5) satisfies
$\|z(t;z_{0})\|\leq\|\exp(At)z_{0}\|+\int_{0}^{t}\|\exp(A(t-\tau))By(\tau)\|d\tau.$
Since $A$ is Hurwitz, $\|\exp(At)z_{0}\|\to 0$ exponentially fast, and the effect of the initial condition $z_{0}$ vanishes from $z(t;z_{0})$ over time $t$. We have $\left\|\exp(At)z_{0}\right\|\leq\operatorname*{cond}(V)e^{\lambda_{\min}(A)t}\|z_{0}\|$, where $\lambda_{\min}(A)$ defined in Section 3.1, and $V$ obtained from the eigendecomposition of $A=V\Lambda V^{-1}$, which is assumed to be diagonalizable. For any $\varepsilon>0$, there exists time $t_{*}$ such that $\left\|\exp(At)z_{0}\right\|\leq\varepsilon$, for all $t\geq t_{*}$, where $t_{*}\coloneqq t_{*}(\varepsilon,z_{0})=\frac{1}{\lambda_{\min}(A)}\ln\left(\frac{\varepsilon}{\operatorname*{cond}(V)\|z_{0}\|}\right).$ Therefore,
$\forall t\geq t_{*},\quad\|z(t;z_{0})\|\leq\varepsilon+\int_{0}^{t}\|\exp(A(t-\tau))By(\tau)\|d\tau.$
@@ 5.2 Data Generation Procedure
As time progresses, the trajectory $z(t;z_{0}^{i})$ becomes almost independent of the initial condition $z_{0}$. This observation leads to the following data generation procedure.
@@
Initial condition sampling: We sample $p\in\mathbb{Z}_{>0}$ initial points $\left\{(x_{0}^{i},z_{0}^{i})\right\}_{i\in[p]}$ uniformly in $\mathcal{X}\times\mathcal{Z}$.
@@
Truncation: For all $i\in[p]$, the samples of the $z$-trajectory for $t_{k}\geq t_{k_{*}}$ are kept and the rest of the samples are discarded, where
@@
This procedure yields the training set
$S_{\mathrm{data}}=\{(x(t_{k};x_{0}^{i}),z(t_{k};z_{0}^{i}))\}_{i\in[p],k\in\{k_{*},\dots,\tau\}}$
containing truncated trajectories with $N_{\mathrm{data}}=p(\tau-k_{*}+1)$ data pairs. We further sample $q\in\mathbb{Z}_{>0}$ points $x_{0}^{j}$ and generate trajectories
$S_{\mathrm{pde}}=\{x(t_{k};x_{0}^{j})\}_{j\in[q],k\in\{1,\dots,\tau\}}$
with $N_{\mathrm{pde}}=q\tau$ collocation points for the PDE. The data $S_{1}=\{S_{\mathrm{data}},S_{\mathrm{pde}}\}$ is used for learning the forward map $\hat{\mathcal{T}}_{\theta}$ with total number of data points $N_{1}=N_{\mathrm{data}}+N_{\mathrm{pde}}$. After training the forward map $\hat{\mathcal{T}}_{\theta}$, its parameters $\theta$ are fixed. We then generate a new dataset $S_{2}=\{(z^{\prime}_{j},x^{\prime}_{j})\}_{j=1}^{N_{2}}$ of size $N_{2}$, where $x^{\prime}$ are randomly sampled points in $\mathcal{X}$ or trajectories from randomly sampled initial points in $\mathcal{X}$, and the corresponding features $z^{\prime}_{j}=\hat{\mathcal{T}}_{\theta}(x^{\prime}_{j})$ are generated from the learned forward map.
Figure 1: Overview of the proposed KKL observer learning algorithm.
@@ 5.3 Empirical Risk Minimization
We approximate the true risks $R_{1}$ and $R_{2}$ in (13) and (15) with their empirical counterparts, computed from finite datasets $S_{1}$ and $S_{2}$. First, we minimize empirical physics-informed risk $\hat{R}_{S_{1}}$ using $S_{1}=(S_{\mathrm{data}},S_{\mathrm{pde}})$ with
$\hat{R}_{S_{1}}(\theta)=\frac{1}{N_{\mathrm{data}}}\sum_{(x,z)\in S_{\mathrm{data}}}\|z-\hat{\mathcal{T}}_{\theta}(x)\|^{2}+\frac{\nu}{N_{\mathrm{pde}}}\sum_{x\in S_{\mathrm{pde}}}\|\mathcal{P}_{\theta}(x)\|^{2}$ (16)
where $\mathcal{P}_{\theta}(x)$ is defined in (12). This risk is the computable objective function for the inverse map and approximates the true risk in (13). Thus, the first-stage empirical risk minimization problem is formulated as
$\min_{\theta\in\mathcal{H}_{\theta}}\hat{R}_{S_{1}}(\theta).$ (17)
We follow the sequential approach, first proposed by [41] for discrete-time systems. After training the forward map $\hat{\mathcal{T}}_{\theta}$, we generate a new dataset $S_{2}=\{(z^{\prime}_{j},x^{\prime}_{j})\}_{j=1}^{N_{2}}$, where $z^{\prime}_{j}=\hat{\mathcal{T}}_{\theta}(x^{\prime}_{j})$. On this dataset $S_{2}$, the empirical reconstruction risk is given by
$\hat{R}_{S_{2}}(\eta\mid\theta)=\frac{1}{N_{2}}\sum_{j=1}^{N_{2}}\|x^{\prime}_{j}-\hat{\mathcal{T}}_{\eta}^{*}(z^{\prime}_{j})\|^{2}$ (18)
which is the computable objective function for the inverse map and approximates the true risk in (15). Thus, the second-stage empirical risk minimization problem is formulated as
$\min_{\eta\in\mathcal{H}_{\eta}}\hat{R}_{S_{2}}(\eta\mid\theta).$ (19)
5.4 Remarks on the Learning Method
@@
Fig. 1 illustrates the workflow of proposed learning algorithm. Our procedure deconstructs the autoencoder framework used in previous works, such as [39], into sequentially learning the maps $\hat{\mathcal{T}}_{\theta}$ and $\hat{\mathcal{T}}_{\eta}^{*}$. This idea serves to address two fundamental issues:
@@
    The joint encoder decoder structure is inherently susceptible to conflicting gradients between components of the objective function, which might result in the optimization getting stuck in bad local minima. This not only degrades the accuracy of the learned solutions but also create the risk of failing to enforce the latent output of the autoencoder to reside within $\mathcal{Z}$, leading to a discrepancy between the input distributions of $\hat{\mathcal{T}}_{\eta}^{*}$ during online operation.